# 11footnotetext: {*} denotes Equal contribution

### v1

@@ 5.2 Wall-clock Time Speedups for Inference
Activation sparsity of a model is not sufficient to directly enable wall-clock time inference speedups (Frantar & Alistarh, 2023). In this subsection, we demonstrate that our custom GPU kernel translates the activation sparsity induced by CATS to real wall-clock time gains.
@@
Figure 3 shows the wall-clock inference time of of the dense model compared to CATS implemented via the custom GPU kernel descripted in Section 4.3, for various sparsity levels of CATS. We evaluate the latency of a single MLP block and the throughput of the generation stage of the end-to-end inference. Mistral-7B contains 32 MLPs with $m=14336$ and $d=4096$, and Llama2-7B contains 32 MLPs with $m=11008$ and $d=4096$ ($m$ and $d$ are defined after Equation 1).
In Figures 3(a) and 3(b), we compare our method (“CATS-with-Custom-Kernel”) with the dense MLP with $m_{dense}=m$ (“Dense”) and the dense MLP with $m_{optimal}=m*Sparsity$ (“Optimal”), the latter of which is a proxy for the best wall-clock time we could hope to achieve. At 50% (respectively, 70%) sparsity, the sparse kernel achieves a $\sim$40(respectively, $\sim$70%) speedup over the original dense MLP. Latency measurements are obtained by doing 20 rounds of warmups, repeating the kernel 80 times, and computing the geometric mean of the latency of each round. The comparison with Dense shows that our sparse kernel can consistently outperform the original MLP. The comparison with Optimal shows that our sparse kernel is close to the Optimal when the sparsity level is low. As the sparsity level increases, the gap between our performance and Optimal increases, as is expected. We note that our sparse kernel performs the same number of memory access as the Optimal but, due to difference in access patterns, the different methods result in different wall-clock time measurements. We note that Optimal can be worse than our sparse kernels when $m_{optimal}$ is not the shape for which GPU libraries have optimized (Tillet & Cox, 2017). Our sparse kernel can be worse than Optimal when the overhead of operations on zero values outweighs the benefit of reduced memory access.
In Figures 4(a) and 4(b), we compare dense models with CATS-with-Custom-Kernel (50% sparsity) on the throughput of the generation stage. The generation stage (or “decoding” stage) is known to be memory-bound (Kim et al., 2023), which suggests CATS can improve inference througput. We test the generation throughput at batch size 1 and beam width 1, and record the latency from the first generated token to the last token. The throughput is calculated by the generated length over latency. The final throughput is averaged (geometric mean) over 50 samples from the RefinedWeb test dataset. CATS can accelerate the generation stage by $\sim$18for Llama2-7B and $\sim$21for Mistral-7B at 50% sparsity.
@@ 6 Discussion and Conclusion
Ethics Statement
We do not foresee any ethical problems with our work as our framework only targets memory and wall-time efficiency of language models.
Reproducibility Statement
We have submitted code to reproduce all of our experiments via a one-line script in our supplementary material; please see the README.md.
Acknowledgements
We thank the anonymous reviewers, area chairs, and program chairs for their reviews and comments of our paper. We will add other acknowledgements for the camera-ready version of this paper.

### v4

@@ 5.3 Wall-clock Time Speedups for Inference
Activation sparsity in a model is not sufficient to directly achieve wall-clock time inference speedups (Frantar & Alistarh, 2023). In this subsection, we demonstrate that our custom GPU kernel translates the activation sparsity induced by CATS into tangible wall-clock time improvements.
@@
Figure 3 shows the wall-clock inference time of the dense model compared to CATS implemented via the custom GPU kernel described in Section 4.3, for various sparsity levels of CATS. We evaluate the latency of a single MLP block and the throughput of the generation stage of the end-to-end inference. Mistral-7B contains 32 MLPs with $m=14336$ and $d=4096$, and Llama2-7B contains 32 MLPs with $m=11008$ and $d=4096$ ($m$ and $d$ are defined after Equation 1).
In Figures 3(a) and 3(b), we compare our method (“CATS-with-Custom-Kernel”) with the dense MLP with $m_{dense}=m$ (“Dense”) and the dense MLP with $m_{optimal}=m*Sparsity$ (“Optimal”), the latter of which is a proxy for the best wall-clock time we could hope to achieve. At 50% and 70% sparsity, the sparse kernel achieves approximately a 40% and 70% speedup, respectively, over the original dense MLP. Latency measurements were obtained by conducting 20 rounds of warmups, repeating the kernel 80 times, and computing the geometric mean of the latency across each round. The comparison with Dense demonstrates that our sparse kernel consistently outperforms the original MLP. The comparison with Optimal shows that our sparse kernel closely approaches the Optimal performance at low sparsity levels. However, as the sparsity level increases, the gap between our performance and Optimal widens, as expected. We note that our sparse kernel performs the same number of memory accesses as the Optimal; however, due to differences in access patterns, the methods result in different wall-clock time measurements. Optimal can perform worse than our sparse kernel when $m_{optimal}$ does not match the shapes optimized by GPU libraries (Tillet & Cox, 2017). Conversely, our sparse kernel can underperform compared to Optimal when the overhead of operations on zero values outweighs the benefits of reduced memory access.
In Figures 4(a) and 4(b), we compare dense models with CATS-with-Custom-Kernel (50% sparsity) on the throughput of the generation stage. The generation stage (or “decoding” stage) is known to be memory-bound (Kim et al., 2023), which suggests CATS can improve inference througput. We test the generation throughput at a batch size of 1 and beam width of 1, and record the latency from the first generated token to the last token. The throughput is calculated by the generated length divided by latency. The final throughput is averaged (geometric mean) over 50 samples from the RefinedWeb test dataset. CATS can accelerate the generation stage by $\sim$18for Llama2-7B and $\sim$21for Mistral-7B at 50% sparsity.
@@ Appendix D Details on Custom GPU Kernel Design
1# PyTorch
2import triton.language as tl
3V = SiLU(X @ W_gate)
4Mask = torch.abs(V) >= threshold
5
6# Triton 1: x_1 = (x @ W_up[Mask]) * v[Mask]
7# Input
8X: [N]
9V: [M]
10W_up: [M,N] # Stored in column-major
11Mask: [M]
12# Output
13X_1: [M]
14# Tunable parameters
15BLOCK_M = {4,8,16,32}
16BLOCK_N = {256,512,1024}
17# Configuration
18grid = (M // BLOCK_M,)
19# Kernel[grid]
20start_m = tl.program_id(0)
21rm = start_m * BLOCK_M + tl.arange(0, BLOCK_M)
22rn = tl.arange(0, BLOCK_N)
23Mask = Mask + rm
24flag = tl.load(Mask) > 0
25W_up = W_up + (rm[:,None] * N + rn[None,:])
26X = X + rn
27acc = tl.zeros((BLOCK_M,), dtype=tl.float32)
28i_mask = flag[:,None]
29for _ in range(N, 0, -BLOCK_N):
30 w = tl.load(W_up, mask=i_mask, other=0.0)
31 x = tl.load(X)
32 acc += tl.sum(w * x[None,:], 1).to(tl.float32)
33 W_up += BLOCK_N
34 X += BLOCK_N
35V = V + rm
36v = tl.load(V, mask=flag, other=0.0)
37acc = acc * v.to(tl.float32)
38X_1 = X_1 + rm
39tl.store(X_1, acc.to(tl.float16), mask=rm < M)
40
41# Triton 2: y = x_1 @ W_down[Mask]
42# Input
43X_1: [M]
44W_down: [M,N]
45Mask: [M]
46# Output
47Y: [N]
48# Tunable parameters
49BLOCK_M = {16,32,64,128}
50BLOCK_N = {128,256,512,1024}
51# Configuration
52grid = (M // BLOCK_M, N // BLOCK_N)
53# Kernel[grid]
54start_m = tl.program_id(0)
55start_n = tl.program_id(1)
56rm = start_m * BLOCK_M + tl.arange(0, BLOCK_M)
57rn = start_n * BLOCK_N + tl.arange(0, BLOCK_N)
58Mask = Mask + rm
59flag = tl.load(Mask) > 0
60W_down = W_down + (rm[:,None] * N + rn[None,:])
61X_1 = X_1 + rm
62w = tl.load(W_down, mask=flag[:,None], other=0.0)
63x = tl.load(X_1)
64acc = tl.sum(a * x[:,None], 0).to(tl.float32)
65Y = Y + rn
66tl.atomic_add(Y, acc.to(tl.float16))
Figure 7: Triton pseudo-code for Algorithm 1.
Appendix E Additional Experiments
We conducted an additional experiment to test whether this trend holds for larger models. We evaluated the Llama 13B model on the same downstream tasks and observed that the performance degradation becomes even more minimal. This leads to the tentative conclusion that CATS becomes more effective for larger models.
Additionally, we compared the performance of CATS with Wanda Sun et al. (2023), a method that prunes weights on a per-output basis based on the product of weight magnitudes and input activation norms. As Wanda achieves unstructured sparsity of weights, it is important to note that this approach may not lead to wall-clock time improvements in practice.
Model WG PIQA SciQ QA HS BoolQ Arc-E Arc-C Avg
Llama2-13B 0.7230 0.7905 0.9460 0.3520 0.6006 0.8055 0.7946 0.4838 0.6870
CATS 50% 0.7111 0.7884 0.9480 0.3540 0.6057 0.7706 0.7870 0.4812 0.6805
CATS 70% 0.6946 0.7862 0.9310 0.3360 0.5987 0.7419 0.7546 0.4497 0.6616
CATS 90% 0.5596 0.6828 0.5820 0.2080 0.3960 0.6416 0.4726 0.3029 0.4807
ReLUfication 0.4933 0.5522 0.2660 0.1440 0.2648 0.4355 0.2757 0.2483 0.3350
Wanda 0.7079 0.7873 0.9450 0.3200 0.5710 0.8122 0.7605 0.4300 0.6667
Table 5: Zero-shot performance of 13B-based models, CATS-based models for varying levels of sparsity, ReLUfication, and Wanda 50% across different benchmark datasets. CATS 50% significantly outperforms other techniques, which suggests CATS scales well with model size.